OpenStudy (anonymous):

Please helP! Find a • b. a = 4i + 8j, b = -2i + 3j

OpenStudy (anonymous):

\[(4i+8j)*(-2i+3j)\] This looks like the dot product of two vectors I'm going to assume

OpenStudy (anonymous):

yes i think so

OpenStudy (anonymous):

multiply the ith components and the jth components with each other so \[4i(-2i)+8j(3j)\] i*i=1 and j*j=1 since these represent unit vectors i*j=0

OpenStudy (anonymous):

\[-8+24=?\]